OpenStudy (anonymous):

Find the general anti- derivative of 5 sec^2 x dx

myininaya (myininaya):

That isn't the general antiderivative but it is a antiderivative. He was looking for a function whose derivative was 5sec^2(x).

OpenStudy (anonymous):

yes

myininaya (myininaya):

Nah loser. She just needs +C. lol

myininaya (myininaya):

There really isn't a lot to do on this problem.

OpenStudy (anonymous):

okay so 5 tan x is correct?

myininaya (myininaya):

Well that is a antiderivative. You are looking for the general antiderivative.

OpenStudy (anonymous):

okay

myininaya (myininaya):

\[\int\limits_{}^{}f(x) dx=F(x)+C , (F'=f)\] C is a constant

myininaya (myininaya):

F(x)+C is the general antiderivative of f(x)

OpenStudy (anonymous):

ok